OpenStudy (anonymous):

can anyone help me integrate x(x^2+1)^3

OpenStudy (anonymous):

help

OpenStudy (anonymous):

another u- sub put \[u=x^2+1\] \[du=2xdx\] so integrate \[\frac{1}{2}\int u^3 du\]

OpenStudy (anonymous):

get \[\frac{u^4}{8}=\frac{(x^2+1)^4}{8}\]

OpenStudy (anonymous):

don't forget the stupid plus C

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

welcomas